### v1

@@ 1 Error-correction: Problem Definition and Key Parameters
The architecture of an error-correcting scheme involves choosing a message space, and an encoding and decoding function. The encoding function maps a message to a codeword over $\Sigma$, where a word is a finite sequence of symbols of $\Sigma$ and codewords are words obtained by encoding some message. This codeword is transmitted over the channel which produces a received word, which may differ from the codeword in at most $p$-fraction of symbols. The receiver now applies the decoding function which maps the received word to a small list (i.e., sequence) of codewords. The transmission is considered successful if the output list contains the message. We note that the classical setting of decoding, which we refer to as unique decoding, required the list size to be $1$, but we will consider the more relaxed version here today.
@@
For simplicity (and ease of comparisons) we will assume the message is a word of length $k$ (for positive integer $k$) over the alphabet $\Sigma$ and so the message space is $\Sigma^{k}$. Thus the encoding function $E:\Sigma^{k}\to\Sigma^{n}$ and a list of $L$ decoder $D:\Sigma^{n}\to(\Sigma^{k})^{L}$. A coding scheme is given by an encoder/decoder pair $(E,D)$ and is considered a $(p,L)$-list decoder if for every $m\in\Sigma^{k}$ and every $p$-bounded channel Ch we have $m\in D(\textsc{Ch}(E(m)))$. (We note that lists conflate the notion of sets and sequences - they are simply meant to be sets, but the decoder outputs the elements of the set in some sequence and this sequence is referred to as a list.) When $L=1$, the $(p,L)$-list decoder is also called a $p$-error-correction scheme.
@@
The broad goal is to study all $5$ parameters jointly, but this can be too much, and so we try to reduce to just two as follows. First we only consider families of $(R,p,q,L)$ coding schemes that can achieve arbitrarily large block lengths. Next we make similar choices for $q$ and $L$, i.e., we allow them to also tend to infinity; however, we have to be careful. First we stress that these are not the only settings of interest. Indeed the most common parameters to study are $q=2$ and $L=1$, but that is not the range of parameters that highlight the role of algebra (though the results do imply something in those settings as well and we may comment on this aspect as side notes below). Also when $q$, $L$ and $n$ are all tending to infinity, one needs to be careful about how they grow relative to each other. For instance allowing $L=q^{n}$ would make coding schemes trivial (as in easy to achieve) and pointless (as in the output list need not give any information about the message being sent). A reasonable choice for the growth of $q$ and $L$ is that each grows as some fixed polynomial in $n$. With these choices the question of attention ends up asking:
@@
Codes of minimum distance $\delta$ also have $(\delta/2,1)$-list decoders and $\delta/2$ is the best possible value here when restricting to $L=1$, for every code of minimum distance $\delta$. when restricted to list size $1$. However allowing larger list sizes sometimes allows us to get better results: In particular, as we will see next, every code of rate $R$ and distance $\delta$ satisfies $R\leq 1-\delta$. Using unique decoding we can thus correct at most (1R)/2(1-R)/2 fraction of errors. But by allowing larger lists we can go all the way to 1R1-R fraction of errors, thus doubling the error-correction capacity at the expense of leaving the receiver with the task of determining which element of the list was the intended message.
@@ 3 The Reed-Solomon Codes
For every function $f:S\to\mathbb{F}_{q}$ with $|S|=n$ there exists a non-zero bivariate polynomial $Q(X,Y)\in\mathbb{F}_{q}[X,Y]$ with $\deg_{X}(Q),\deg_{Y}(Q)\leq\sqrt{n}$ such that for all $a\in S$, $Q(a,r(a))=0$.
@@
On the one hand the lemma above is “trivial” — it made no assumptions on the function $r$ to prove the existence of $Q$. On the other hand it is algebraically quite powerful since the algebraic degree of $Q$ is just twice $\sqrt{n}$. To see why it helps with list decoding, consider a setting where $t>2k\sqrt{n}$ and the transmitted codeword is $p(X)$. In this case the polynomials $Q(X,Y)$ and R(X,Y)=Yp(X)R(X,Y)=Y-p(X) share $t$ common zeroes in the plane $\mathbb{F}_{q}\times\mathbb{F}_{q}$. Specifically every point $a\in S$ such that $r(a)=p(a)$ satisfies $Q(a,r(a))=R(a,r(a))=0$. A classical result from the origins of algebraic geometry, namely Bézout’s theorem in the plane, tells us that polynomials $Q$ and $R$ of degree $d_{1}$ and $d_{2}$ respectively can share more that $d_{1}d_{2}$ common zeroes only if they share a common factor. In our setting $d_{1}=2\sqrt{n}$ and $d_{2}=k$ and so if $t>2k\sqrt{n}$ then it follows that $Q(X,Y)$ and Yp(X)Y-p(X) share a common factor. But the latter polynomial is irreducible and so it must be an irreducible factor of $Q(X,Y)$. And this is useful since a bivariate polynomial can be factored into irreducible factors in polynomial time over finite fields (and many other settings) due to an illustrious line of work including [Ber70, LLL82, Kal85, Len86, Gri84, vzGG13].
@@
The correctness follows from the fact, by Lemma 4.1 a polynomial $Q$ satisfying the conditions of Step 1 always exists (and so Step 1 will return such a polynomial); and by Lemma 4.2 any polynomial satisfying the conditions of Step 1 will have yp(X)y-p(X) as an irreducible factor, provided $t>D_{x}+D_{y}\cdot d$ where $D_{x}=D_{x}=\sqrt{n}$ and d=k1d=k-1. ∎
@@ 5 Optimizing the List-decoding algorithm: Weighted-degrees and Multiplicities
But first, let’s mention what is known about the limits of $t$ for which the problem may be solvable in polynomial time. Note that to output a list of size $L$ the algorithm must run in time at least $L$ and so the question we focus on is: “For what value of $t$ is $L$ known to be polynomially bounded in nn?”. When $t\geq\frac{n+k}{2}$ we are at the unique decoding limit, i.e., here we know $L=1$. When $t>\sqrt{2kn}$ an inclusion-exclusion argument [^4] Suppose there are $L$ polynomials $p_{1},\ldots,p_{L}$ that agree with the received word on $t$ points, then we have that there at $L$ sets $S_{1},\ldots,S_{L}$ such $S_{i}\subseteq S$ with $|S_{i}|\geq t$ and $|S_{i}\cap S_{j}|\leq k-1$ for $i\neq j$. Using an “inclusion-exclusion argument”, i.e., the sequence of inequalities $n\geq|\cup_{i}S_{i}|\geq\sum_{i}|S_{i}|-\sum_{i<j}|S_{i}\cap S_{j}|\geq Lt-\binom{L+1}{2}(k-1)$ one can conclude that $L\leq 2n/k$ if $t\geq\sqrt{2nk}$. In particular we get $L$ is polynomially bounded in $n$ (and even a constant if $n=O(k)$. This new bound is better than (smaller than) the previous one when $k$ is relatively small compared to $n$, but for say $k=n/2$ the previous bound was better. The final improvement, known as the Johnson bound, shows that when $t>\sqrt{kn}$ then $L$ grows polynomially in $n$. When $L$ is larger we really don’t know the answer: There are results suggesting that $t\leq\sqrt{kn}$ could really lead to superpolynomial list sizes for some sets $S\subseteq\mathbb{F}_{q}$ over special fields $\mathbb{F}_{q}$ (notably the work of Ben-Sasson, Kopparty and Radhakrishnan [BSKR06]). Other recent results show that for random sets $S\subseteq\mathbb{F}_{q}$ of sufficiently small size, one could imagine going down to $t=k+o(n)$ (see for instance, Alrabiah, Guruswami and Li [AGL24] and references therein). So beyond the Johnson bound, the answer seems open. In the rest of this section we will show how to achieve the Johnson bound efficiently, i.e., with a polynomial time decoder.
@@ Acknowledgments
I would like to thank Kumar Murty for encouraging me to write this article. The article benefits from discussions over the years with my colleagues and collaborators and I’d like to thank Siddharth Bhandari, Oded Goldreich, Venkatesan Guruswami, Prahladh Harsha, Swastik Kopparty, Mrinal Kumar, Atri Rudra, Ramprasad Saptharishi, Shubhangi Saraf, Salil Vadhan, and Avi Wigderson for improving my understanding of the results covered here. Thanks to Prashanth Amireddy, Cassandra Marcussen and Aaron (Louie) Putterman for proofreading this article and catching many errors. Thanks to an anonymous reviewer for their careful review that caught several errors and point out inconsistencies.

### v2

@@ 1 Error-correction: Problem Definition and Key Parameters
The architecture of an error-correcting scheme involves choosing a message space, and an encoding and decoding function. The encoding function maps a message to a codeword over $\Sigma$, where a word is a finite sequence of symbols of $\Sigma$ and codewords are words obtained by encoding some message. This codeword is transmitted over the channel which produces a received word, which may differ from the codeword in at most $p$-fraction of symbols. The receiver now applies the decoding function which maps the received word to a small list (i.e., sequence) of codewords. The transmission is considered successful if the output list contains the message. We note that the classical setting of decoding, which we refer to as unique decoding, required the list size to be $1$, but we will consider the more relaxed version in this article.
@@
For simplicity (and ease of comparisons) we will assume the message is a word of length $k$ (for positive integer $k$) over the alphabet $\Sigma$ and so the message space is $\Sigma^{k}$. Thus the encoding is a function $E:\Sigma^{k}\to\Sigma^{n}$ and a list of $L$ decoder is a function $D:\Sigma^{n}\to(\Sigma^{k})^{L}$. A coding scheme is given by an encoder/decoder pair $(E,D)$ and is considered a $(p,L)$-list decoder if for every $m\in\Sigma^{k}$ and every $p$-bounded channel Ch we have $m\in D(\textsc{Ch}(E(m)))$. (We note that lists conflate the notion of sets and sequences - they are simply meant to be sets, but the decoder outputs the elements of the set in some sequence and this sequence is referred to as a list.) When $L=1$, the $(p,L)$-list decoder is also called a $p$-error-correction scheme.
@@
The broad goal is to study all $5$ parameters jointly, but this can be too much, and so we try to reduce to just two as follows. First we only consider families of $(R,p,q,L)$ coding schemes that can achieve arbitrarily large block lengths. Next we make similar choices for $q$ and $L$, i.e., we allow them to also tend to infinity; however, we have to be careful. First we stress that these are not the only settings of interest. Indeed the most common parameters to study are $q=2$ and $L=1$, but that is not the range of parameters that highlight the role of algebra (though the results do imply something in those settings as well and we may comment on this aspect as side notes below). Also when $q$, $L$ and $n$ are all tending to infinity, one needs to be careful about how they grow relative to each other. For instance allowing $L=q^{k}$ would make coding schemes trivial (as in easy to achieve) and pointless (as in the output list need not give any information about the message being sent). A reasonable choice for the growth of $q$ and $L$ is that each grows as some fixed polynomial in $n$. With these choices the question of attention ends up asking:
@@
Codes of minimum distance $\delta$ also have $(\delta/2,1)$-list decoders. Conversely for every code of minimum distance $\delta$, $\delta/2$ is the best possible value for $\rho$ when restricted to list size $L=1$. However allowing larger list sizes sometimes allows us to get better results: In particular, as we will see next, every code of rate $R$ and distance $\delta$ satisfies $R\leq 1-\delta$. Using unique decoding we can thus correct at most (1R)/2(1-R)/2 fraction of errors. But by allowing larger lists we can go all the way to 1R1-R fraction of errors, thus doubling the error-correction capacity at the expense of leaving the receiver with the task of determining which element of the list was the intended message.
@@ 3 The Reed-Solomon Codes
For every function $r:S\to\mathbb{F}_{q}$ with $|S|=n$ there exists a non-zero bivariate polynomial $Q(X,Y)\in\mathbb{F}_{q}[X,Y]$ with $\deg_{X}(Q),\deg_{Y}(Q)\leq\sqrt{n}$ such that for all $a\in S$, $Q(a,r(a))=0$.
@@
On the one hand the lemma above is “trivial” — it made no assumptions on the function $r$ to prove the existence of $Q$. On the other hand it is algebraically quite powerful since the algebraic degree of $Q$ is just twice $\sqrt{n}$. To see why it helps with list decoding, consider a setting where $t>2k\sqrt{n}$ and the codeword that the sender sends (before errors) is $p(X)$. In this case the polynomials $Q(X,Y)$ and R(X,Y)=Yp(X)R(X,Y)=Y-p(X) share $t$ common zeroes in the plane $\mathbb{F}_{q}\times\mathbb{F}_{q}$. Specifically every point $a\in S$ such that $r(a)=p(a)$ satisfies $Q(a,r(a))=0=R(a,r(a))$, where the former equality holds for all $a$ and the latter holds since $r(a)=p(a)$. A classical result from the origins of algebraic geometry, namely Bézout’s theorem in the plane, tells us that polynomials $Q$ and $R$ of degree $d_{1}$ and $d_{2}$ respectively can share more that $d_{1}d_{2}$ common zeroes only if they share a common factor. In our setting $d_{1}=2\sqrt{n}$ and $d_{2}=k$ and so if $t>2k\sqrt{n}$ then it follows that $Q(X,Y)$ and Yp(X)Y-p(X) share a common factor. But the latter polynomial is irreducible and so it must be an irreducible factor of $Q(X,Y)$. And this is algorithmically effective since a bivariate polynomial can be factored into irreducible factors in polynomial time over finite fields (and many other settings) due to an illustrious line of work including [Ber70, LLL82, Kal85, Len86, Gri84, vzGG13].
@@
The correctness follows from the fact, by Lemma 4.1 a polynomial $Q$ satisfying the conditions of Step 1 always exists (and so Step 1 will return such a polynomial); and by Lemma 4.2 any polynomial satisfying the conditions of Step 1 will have yp(X)y-p(X) as an irreducible factor, provided $t>D_{x}+D_{y}\cdot d$ where $D_{x}=D_{y}=\sqrt{n}$ and d=k1d=k-1. ∎
@@ 5 Optimizing the List-decoding algorithm: Weighted-degrees and Multiplicities
But first, let’s mention what is known about the limits of $t$ for which the problem may be solvable in polynomial time. Note that to output a list of size $L$ the algorithm must run in time at least $L$ and so the question we focus on is: “For what value of $t$ is $L$ known to be polynomially bounded in nn?”. When $t\geq\frac{n+k}{2}$ we are at the unique decoding limit, i.e., here we know $L=1$. When $t>\sqrt{2kn}$ an inclusion-exclusion argument shows that $L\leq 2n/k$ if $t\geq\sqrt{2nk}$ [^4] Suppose there are $L$ polynomials $p_{1},\ldots,p_{L}$ that agree with the received word on $t$ points, then we have that there at $L$ sets $S_{1},\ldots,S_{L}$ such $S_{i}\subseteq S$ with $|S_{i}|\geq t$ and $|S_{i}\cap S_{j}|\leq k-1$ for $i\neq j$. Using an “inclusion-exclusion argument”, i.e., the sequence of inequalities $n\geq|\cup_{i}S_{i}|\geq\sum_{i}|S_{i}|-\sum_{i<j}|S_{i}\cap S_{j}|\geq Lt-\binom{L+1}{2}(k-1)$ one can conclude that $L\leq 2n/k$ if $t\geq\sqrt{2nk}$. In particular we get $L$ is polynomially bounded in $n$ (and even a constant if $n=O(k)$.. This new bound is better than (smaller than) the previous one when $k$ is relatively small compared to $n$, but for say $k=n/2$ the previous bound was better. The final improvement, known as the Johnson bound, shows that when $t>\sqrt{kn}$ then $L$ grows polynomially in $n$. When $L$ is larger we really don’t know the answer: There are results suggesting that $t\leq\sqrt{kn}$ could really lead to superpolynomial list sizes for some sets $S\subseteq\mathbb{F}_{q}$ over special fields $\mathbb{F}_{q}$ (notably the work of Ben-Sasson, Kopparty and Radhakrishnan [BSKR06]). Other recent results show that for random sets $S\subseteq\mathbb{F}_{q}$ of sufficiently small size, one could imagine going down to $t=k+o(n)$ (see for instance, Alrabiah, Guruswami and Li [AGL24] and references therein). So beyond the Johnson bound, the answer seems open. In the rest of this section we will show how to achieve the Johnson bound efficiently, i.e., with a polynomial time decoder.
@@ Acknowledgments
I would like to thank Kumar Murty for encouraging me to write this article. The article benefits from discussions over the years with my colleagues and collaborators and I’d like to thank Siddharth Bhandari, Oded Goldreich, Venkatesan Guruswami, Prahladh Harsha, Swastik Kopparty, Mrinal Kumar, Atri Rudra, Ramprasad Saptharishi, Shubhangi Saraf, Srikanth Srinivasan, Salil Vadhan, and Avi Wigderson for improving my understanding of the results covered here. Thanks to Prashanth Amireddy, Oded Goldreich, Cassandra Marcussen and Aaron (Louie) Putterman for proofreading this article and catching many errors. Thanks to an anonymous reviewer for their careful review that caught several errors and point out inconsistencies.